# Time-domain study of coupled collective excitations in quantum materials

### v1

@@ Abstract
Quantum materials hold immense promises for future applications due to their intriguing electronic, magnetic, thermal, and mechanical properties that are often traced to a complex interplay among different microscopic degrees of freedom. Important insights of such interactions come from studying the collective excitations of electrons, spins, orbitals, and lattice, whose cooperative motions play a crucial role in determining the novel behavior of these systems and offer us a key tuning knob to modify material properties on-demand through external perturbations. In this regard, ultrafast light-matter interaction has shown great potential in controlling the couplings of collective excitations, and rapid progress in a plethora of time-resolved techniques down to the attosecond regime has significantly advanced our understanding of the coupling mechanism and guided us in manipulating the dynamical property of quantum materials. This review aims to highlight recent experiments on visualizing collective excitations in the time domain, focusing on the coupling mechanism between different collective modes such as phonon-phonon, phonon-magnon, phonon-exciton, magnon-magnon, magnon-exciton, and various polaritons. We introduce how these collective modes are excited by an ultrashort laser pulse and probed by different ultrafast techniques, and we explain how the coupling between collective excitations governs the ensuing nonequilibrium dynamics. We also provide some perspectives on future studies that can lead to further discoveries regarding the emergent properties of quantum materials both in and out of equilibrium.
@@ I Introduction
Quantum materials have garnered considerable interest due to their remarkable properties associated with, for instance, unconventional superconductivity 1, 2, 3, 4, multiferroicity 5, colossal magnetoresistance 6, and topological quantum states 7. These properties hold great promise for future applications 8 such as topological quantum computing 9 and spintronic devices 10, yet a precise understanding of the underlying mechanism remains elusive in many cases. At the microscopic level, these properties emerge as a result of complex interactions among the charge, spin, orbital, and lattice degrees of freedom. Take superconductivity as an example: the formation of Cooper pairs in the ground state is either aided by electron-phonon interaction in conventional superconductors 11, 12 or other types of electron-boson coupling, if any, in unconventional superconductors. Unraveling the interplay among these degrees of freedom is crucial for advancing our knowledge of the rich phase diagrams of quantum systems. However, the key challenge has been the coexistence of multi-way couplings of a myriad of quasiparticles and excitations, which makes it difficult to isolate a specific type of interaction we are interested in.
@@
To tackle this challenge, recent advances in time-resolved techniques based on ultrafast light-matter interaction have identified time as a new dimension to investigate the couplings among the microscopic degrees of freedom. Leveraging the time resolution of ultrashort pulses down to the attosecond level, experiments based on the pump-probe protocol offer a unique way to disentangle the intertwined couplings at the intrinsic timescales of microscopic interactions 13, 14. Aside from studying the dynamics of individual quasiparticles, a cooperative and wave-like motion of different degrees of freedom — referred to as collective excitations — can be directly observed in the time domain. One of the most notable examples is photoinduced coherent phonons 15, 16, where a macroscopic number of atoms move in perfect synchronization. Such lattice dynamics show up in a variety of time-resolved probes, such as the periodic oscillations of diffraction peak intensity in time-resolved scattering 17, 18, 19, 20, 21, 22 or the oscillation of optical reflectivity and transmissivity in transient optical spectroscopy 23, 24, 25. Oftentimes, more than one collective excitations emerge after a photoexcitation event, and their coupling is governed by the underlying interaction among the constituent degrees of freedom. Therefore, investigating these collective modes and their couplings provides an important perspective to study the microscopic interactions in quantum materials, which in turn offers an opportunity to use light to modify dynamical material properties.
@@
This review focuses on recent experimental progress in understanding the couplings between collective excitations thanks to the development of ultrafast techniques with ever better spatiotemporal resolution and the capability to generate intense laser pulses with highly tailored properties. We begin in Sec. II by explaining why we are interested in studying collective excitations and how time-domain probes of these excitations and their couplings bring out new insights into material properties that are otherwise missing in equilibrium probes. Given that a single probe cannot capture all the degrees of freedom and their collective modes, we subsequently introduce some select ultrafast techniques in Sec. III. In Sec. IV, we delve deeper into the couplings of collective excitations by introducing specific cases such as photon-phonon, phonon-phonon, phonon-magnon, phonon-exciton, among other coupling mechanisms. An outlook for future studies in this field is included in Sec. V.
@@
As this review does not aim to cover the more general area of ultrafast science in condensed matter systems yet an understanding of the broader field is helpful for contextualizing the present article, we hereby refer readers to a more general survey of theoretical advances and experimental techniques in ref. 26 as well as a few other topical reviews. For techniques, see the following review articles on ultrafast optical spectroscopy 27, 28, 29, 30, 31, 25, 32, ultrafast electron diffraction (UED) and microscopy 33, 34, 35, time-resolved X-ray scattering 36, 37, 38, 39, 35, and time- and angle-resolved photoemission spectroscopy (tr-ARPES) 40, 41, 42, 43, 44. Regarding precision engineering of material properties through ultrafast light pulses, we refer readers to excellent reviews on nonequilibrium engineering such as nonlinear phononics 45, 46, 47 and Floquet engineering 48, 49. From the perspective of different material classes, readers can check out reviews on two-dimensional materials 50, magnetic materials 51, high-temperature superconductors 30, 46, complex oxides 52, and topological insulators 49. These lists are by no means exhaustive but provide sufficient conceptual background for appreciating the experiments highlighted in this review.
@@ II.1 Collective excitations in and out of equilibrium
In equilibrium, collective excitations can reveal crucial information about the ground state orders 57 and the associated phase transitions due to a symmetry change. For example, structural or magnetic transitions are often accompanied by soft phonons or magnons, respectively, which signify an instability in the system. In addition, these collective modes are highly responsive to external stimuli, making it possible to achieve on-demand properties of quantum materials as one applies electromagnetic fields, pressure, strain, or ultrafast photoexcitation to modify the collective dynamics 58. While collective excitations are readily detected in an equilibrium setting and have hallmarks in a variety of measurements ranging from transport to scattering experiments, the underlying microscopic motions are incoherent and hence only time-averaged dynamics are obtained.
This situation changed with the development of ultrafast techniques that are capable of visualizing microscopic motion with a time resolution commensurate with the characteristic period of collective excitations. Under the paradigm of pump-probe measurements, coherent dynamics can be excited by an ultrashort laser pulse, offering a general framework for inducing and controlling coherent dynamics in solids. For example, coherent phonons or magnons are often observed following a photoexcitation event, where ionic displacement or spin precession occurs in complete synchrony with the same phase over a macroscopic sample volume. Such coherent motions are in stark contrast to the random movements that account for incoherent phonons and magnons in thermal equilibrium.
@@ II.2 Photoinduced phase transitions
The ability to observe these coherent dynamics yields critical information about light-induced phase transitions. One extensively studied example is transitions that involves a symmetry change under the Ginzburg-Landau framework, such as those found in material systems hosting a charge density wave (CDW). One type of collective excitations of the CDW ground state is the amplitude mode, which has been widely observed following above-gap photoexcitation that transiently suppresses the order parameter amplitude and hence launches a coherent amplitude oscillation 59, 60, 61. Importantly, the photo-driven amplitude mode, in particular its magnitude and frequency, can inform us how the phase transition unfolds after the laser pulse incidence, such as whether the phase change proceeds in a spatially homogeneous fashion or heterogeneity like topological defects is involved 62, 63. As another example, in a prototypical CDW material LaTe3, the CDW order parameter can be suddenly quenched within a time set by the characteristic time of the amplitude mode oscillations 23. With varying photoexcitation densities, the suppression of the CDW order takes the longest time at the threshold photoexcitation density at which the order parameter transiently vanishes, demonstrating dynamical slowing-down in a nonequilibrium transition. Using more than one pump pulses, one can further harness the coherence in photoinduced collective excitations to precisely control the extent of the nonequilibrium transition. This coherent control was recently demonstrated in 1D Si(111)-In nanowires 64 and quasi-2D 1-TaS2 (ref. 65), both of which are archetypical CDW systems.
While most collective modes studied in the context of photoinduced phase transitions are driven by above-gap photoexcitation with ultraviolet to near-infrared light, they can also be induced by an intense mid-infrared or terahertz field. Instead of photo-doping and injecting mobile carriers into the solids, long-wavelength light can directly excite (i) dipole-active modes through the oscillating electric field, (ii) Raman-active modes through two-photon processes 66, or (iii) other strong-field effects 67 to initiate the collective mode that subsequently drives a transition. Novel metastable states and phenomenology can emerge, such as dynamic multiferroicity 68, 69, 70, 71, a change in electronic topology 72, 73, and magnetism induced by light-driven chiral phonons 74, 75, 76 or as a result of strong spin-lattice coupling 77. We again take CDW materials as an example. In metallic 3-Ta1+xSe2, when its CDW amplitude mode is excited by an intense terahertz pulse, an energy gap appears in the optical conductivity spectrum and executes a dynamical evolution in sync with the amplitude mode oscillation, suggesting a transition into a transient insulating state that has no equilibrium counterpart 78.
@@ II.3 Dressed states
Aside from the role of collective modes during the course of a nonequilibrium phase transition, they can also help induce novel dynamical states of matter via coherent light-matter interactions. For instance, whereas Floquet-Bloch states have been conventionally realized by hybridization between electronic states and an external periodic light field 79, 80, 81, 82, 83, 84, 85, there are a number of theoretical proposals for the Floquet states achieved by replacing photons with phonons 86, 87 or excitons 88. One signature of Floquet-Bloch states is the appearance of electronic band replicas detected in tr-ARPES experiments, where replica bands follow specific intensity relations and polarization selection rules of the exciting photons. Importantly, these bands are spectrally spaced apart by , where is the angular frequency of the bosonic field, be it a photon, phonon, or exciton. In a recent photoemission experiments using a momentum microscope that was specifically designed to study very small samples, exciton-induced Floquet effects were reported in monolayer WS2 in the presence of a high density of resonantly-driven photoexcited excitons 89, which also bear the hallmark of an excitonic insulator phase.
@@ II.4 Phase resolution of coherent collective excitations
When coherent dynamics are launched in solids upon photoexcitation, the motion of collective excitations manifests as wave-like oscillations. The waveform of these oscillations can be directly visualized via an ultrafast probe, an example of which is shown in Fig. 1b that depicts an oscillating Kerr rotation signal caused by a coherent magnon. Therefore, the phase of the collective excitation oscillations can be determined (labeled as in Fig. 1b, bottom panel), which is important information that equilibrium experiments cannot measure directly.
@@
Phase-sensitivity can be used to isolate a single collective mode that affects multiple degrees of freedom, providing a way to measure the coupling strength between them 19, 22. More specifically, by selectively inducing a particular oscillatory response using a tailored pump, and by leveraging phase-locked coherent features from different observables for different degrees of freedom, one can obtain the coupling strength directly, independent of the theoretical model. For example, following the excitation of a coherent phonon, tr-ARPES can probe spectral shifts in the electronic band energy while time-resolved X-ray diffraction (tr-XRD) can quantify the shifts in each atomic position for the phonon of interest. By combining these two measurements, electron-phonon coupling of a specific electronic band to a specific phonon mode can be calculated. This procedure has been applied to an iron-based superconductor FeSe (ref. 19) and three-dimensional topological insulators Bi2Se3 and Bi2Te3 (ref. 22). The extracted electron-phonon coupling strength in Bi2Se3 and Bi2Te3 generally conforms to the expectation from density functional theory (DFT) computations while that in FeSe shows a significant deviation from DFT, which underestimates the strength. Instead, after electron-electron correlation effects were included in a self-consistent DFT-dynamical mean field theory, a quantitative agreement emerged, highlighting an enhancement of the coupling strength due to strong electron correlations. Besides electron-phonon coupling, this multi-probe experimental protocol has also been deployed to measure the dynamical magnetoelectric coupling in multiferroic materials. For instance, in NiI2, a giant coupling strength was obtained 90 where a single electromagnon mode induces periodic modulations in both the dipolar and magnetic orders in NiI2, revealing an interesting /2 phase shift and lock-in behavior between electric polarization and magnetization oscillations that are at terahertz frequencies.
@@
Even with a single experimental probe, a changing oscillation phase with respect to some tuning parameter can be used to reveal material properties 91, 92, 93 and to identify microscopic processes underlying the collective excitation 94. For instance, far above the Curie temperature, the optical phonon mode in a layered ferromagnetic insulator CrSiTe3 can be excited due to the conventional displacive mechanism 15, but the coherent phonon phase experiences a shift when the material cools below a temperature scale where in-plane ferromagnetic correlations start to develop 92. Such a phase shift was only observed when the excitation photon energy is sufficiently large to induce a charge transfer from the Te2- to Cr3+ ions, a process that also increases the superexchange energy. The phase shift was interpreted as evidence for a coherent phonon mechanism due to superexchange enhancement via magneto-elastic coupling, which only applies in the presence of short-range ferromagnetic correlation in this material. Notably, the phase shift of an mode across the Néel temperature also appears in an antiferromagnetic insulator FePS3, providing evidence for a selective coupling between the magnetic order to the trigonal distortions in this layered system 93. In another example of a coherent acoustic phonon in a freestanding film of La2/3Ca1/3MnO3, three different observables in time-resolved diffraction (Bragg peak intensity, width, and position) were observed to synchronously oscillate with a particular phase relation — either in-phase or out-of-phase — after photoexcitation 94. By theoretically modeling the phase relation across the distinct observables from this coherent acoustic mode, Su and coworkers interpreted it as features of ultrafast wrinkle formation that results from delamination between the freestanding film and the substrate at its boundary, offering unique insights into the oft-neglected film–substrate interaction that governs the properties of freestanding structures.

### v2

@@ Abstract
Quantum materials hold immense promises for future applications due to their intriguing electronic, magnetic, thermal, and mechanical properties that often arise from a complex interplay between microscopic degrees of freedom. Important insights of such interactions come from studying the collective excitations of electrons, spins, orbitals, and lattice, whose cooperative motions play a crucial role in determining the novel behavior of these systems and offer us a key tuning knob to modify material properties on-demand through external perturbations. In this regard, ultrafast light-matter interaction has shown great potential in controlling the couplings of collective excitations, and rapid progress in a plethora of time-resolved techniques down to the attosecond regime has significantly advanced our understanding of the coupling mechanisms and guided us in manipulating the dynamical properties of quantum materials. This review aims to highlight recent experiments on visualizing collective excitations in the time domain, focusing on the coupling mechanisms between different collective modes such as phonon-phonon, phonon-magnon, phonon-exciton, magnon-magnon, magnon-exciton, and various polaritons. We introduce how these collective modes are excited by an ultrashort laser pulse and probed by different ultrafast techniques, and we explain how the coupling between collective excitations governs the ensuing nonequilibrium dynamics. We also provide some perspectives on future studies that can lead to discoveries of the emergent properties of quantum materials both in and out of equilibrium.
@@ I Introduction
Quantum materials have garnered considerable interest due to their remarkable properties associated with, for instance, unconventional superconductivity 1, 2, 3, 4, multiferroicity 5, colossal magnetoresistance 6, and topological quantum states 7. These properties hold great promise for future applications 8 such as topological quantum computing 9 and spintronic devices 10, yet a precise understanding of the underlying mechanisms remains elusive in many cases. At the microscopic level, these properties emerge as a result of complex interactions among the charge, spin, orbital, and lattice degrees of freedom. Take superconductivity as an example: the formation of Cooper pairs in the ground state is either aided by electron-phonon interaction in conventional superconductors 11, 12 or other types of electron-boson coupling, if any, in unconventional superconductors. Unraveling the interplay among these degrees of freedom is crucial for advancing our knowledge of the rich phase diagrams of quantum systems. However, one key challenge has been the coexistence of multi-way couplings of a myriad of quasiparticles and excitations, which makes it difficult to isolate a specific type of interaction we are interested in.
@@
To tackle this challenge, recent advances in time-resolved techniques based on ultrafast light-matter interaction have identified time as a new dimension to investigate the couplings among the microscopic degrees of freedom. Leveraging the time resolution of ultrashort pulses down to the attosecond level, experiments based on the pump-probe protocol offer a unique way to disentangle the intertwined couplings at the intrinsic timescales of microscopic interactions 13, 14. Aside from studying the dynamics of individual quasiparticles, a cooperative and wave-like motion of different degrees of freedom — referred to as collective excitations — can be directly observed in the time domain. One of the most notable examples is photoinduced coherent phonons 15, 16, where a macroscopic number of atoms move in perfect synchronization. Such lattice dynamics show up in a variety of time-resolved probes, such as the periodic oscillations of diffraction peak intensity in time-resolved scattering 17, 18, 19, 20, 21, 22 and the oscillations of optical reflectivity and transmissivity in transient optical spectroscopy 23, 24, 25. Oftentimes, more than one type of collective excitation emerges after a photoexcitation event, and their coupling is governed by the underlying interaction among the constituent degrees of freedom. Investigating these collective modes and their couplings provides an important perspective to study the microscopic interactions in quantum materials, which in turn offers an opportunity to use light to modify dynamical material properties.
@@
This review focuses on recent experimental progress in understanding the couplings between collective excitations thanks to the development of ultrafast techniques with ever better spatiotemporal resolution and the capability to generate intense laser pulses with highly tailored properties. We begin in Sec. II by explaining why we are interested in studying collective excitations and how time-domain probes of these excitations and their couplings bring out new insights into material properties that are otherwise missing in equilibrium probes. Given that a single probe cannot capture all the degrees of freedom and their collective modes, we subsequently introduce some select ultrafast techniques in Sec. III. In Sec. IV, we delve deeper into the couplings of collective excitations by introducing specific cases such as photon-phonon, phonon-phonon, phonon-magnon, phonon-exciton, and other coupling mechanisms. An outlook for future studies in this field is included in Sec. V.
@@
While this review does not aim to cover the more general area of ultrafast science in condensed matter systems, such understanding of the broader field is helpful for contextualizing the present article. Therefore, we refer readers to a more general survey of theoretical advances and experimental techniques in ref. 26 as well as a few other topical reviews. For techniques, see the following review articles on ultrafast optical spectroscopy 27, 28, 29, 30, 31, 25, 32, ultrafast electron diffraction (UED) and microscopy 33, 34, 35, time-resolved X-ray scattering 36, 37, 38, 39, 35, and time- and angle-resolved photoemission spectroscopy (tr-ARPES) 40, 41, 42, 43, 44. Regarding precision engineering of material properties through ultrafast light pulses, we refer readers to excellent reviews on nonequilibrium engineering such as nonlinear phononics 45, 46, 47 and Floquet engineering 48, 49. From the perspective of different material classes, see reviews on two-dimensional materials 50, magnetic materials 51, high-temperature superconductors 30, 46, complex oxides 52, and topological insulators 49. These lists are by no means exhaustive but provide sufficient conceptual background for appreciating the experiments highlighted in this review.
@@ II.1 Collective excitations in and out of equilibrium
In equilibrium, collective excitations can reveal crucial information about the ground state orders 57 and the associated phase transitions due to a symmetry change. For example, structural or magnetic transitions are often accompanied by soft phonons or magnons, respectively, which signify an instability in the system. In addition, these collective modes are highly responsive to external stimuli, making it possible to achieve on-demand properties of quantum materials as one applies electric or magnetic fields, pressure, strain, or ultrafast photoexcitation to modify the collective dynamics 58. While collective excitations are readily detected in an equilibrium setting and have hallmarks in a variety of measurements ranging from transport to scattering experiments, the underlying microscopic motions are incoherent and hence only time-averaged dynamics are obtained.
This situation changed with the development of ultrafast techniques that are capable of visualizing microscopic motion with a time resolution commensurate with the characteristic period of collective excitations. Under the paradigm of pump-probe measurements, coherent dynamics can be excited by an ultrashort laser pulse, offering a general framework for inducing and controlling coherent dynamics in solids. For example, coherent phonons or magnons are often observed following a photoexcitation event, where ionic displacement or spin precession occurs in complete synchrony with the same temporal phase over a macroscopic sample volume. Such coherent motions are in stark contrast to the random movements that account for incoherent phonons and magnons in thermal equilibrium.
@@ II.2 Photoinduced phase transitions
The ability to observe these coherent dynamics yields critical information about light-induced phase transitions. One extensively studied example is transitions that involves a symmetry change under the Ginzburg-Landau framework, such as those found in material systems hosting a charge density wave (CDW). One type of collective excitations of the CDW ground state is the amplitude mode, which has been widely observed following an above-gap photoexcitation that transiently suppresses the order parameter amplitude and hence launches a coherent amplitude oscillation 59, 60, 61. Importantly, the photo-driven amplitude mode, in particular its magnitude and frequency, can inform us how the phase transition unfolds after the laser pulse incidence, such as whether the phase change proceeds in a spatially homogeneous fashion or involves heterogeneity like the formation of topological defects62, 63. As another example, in a prototypical CDW material LaTe3, the CDW order parameter can be suddenly quenched within a time set by the characteristic time of the amplitude mode oscillations 23. With varying photoexcitation densities, the suppression of the CDW order takes the longest time at the threshold photoexcitation density at which the order parameter transiently vanishes, demonstrating dynamical slowing-down in a nonequilibrium transition. Using more than one pump pulses, one can further harness the coherence in photoinduced collective excitations to precisely control the extent of the nonequilibrium transition. This coherent control was recently demonstrated in 1D Si(111)-In nanowires 64 and quasi-2D 1-TaS2 (ref. 65), both of which are archetypical CDW systems.
While most collective modes studied in the context of photoinduced phase transitions are driven by above-gap photoexcitation with ultraviolet to near-infrared light, they can also be induced by an intense mid-infrared or terahertz field. Instead of photo-doping and injecting mobile carriers into the solids, long-wavelength light can directly excite (i) dipole-active modes through the oscillating electric field, (ii) Raman-active modes through two-photon processes 66, or (iii) other strong-field effects 67 to initiate the collective mode that subsequently drives a transition. Novel metastable states and phenomenology can emerge, such as dynamic multiferroicity 68, 69, 70, 71, a change in electronic topology 72, 73, and magnetism induced by light-driven chiral phonons 74, 75, 76 or as a result of strong spin-lattice coupling 77. We again use a CDW material as an example. In metallic 3-Ta1+xSe2, when its CDW amplitude mode is excited by an intense terahertz pulse, an energy gap appears in the optical conductivity spectrum and executes a dynamical evolution in sync with the amplitude mode oscillation, suggesting a transition into a transient insulating state that has no equilibrium counterpart 78.
@@ II.3 Dressed states
Collective modes, aside from their role during the course of a nonequilibrium phase transition, can also help induce novel dynamical states of matter via coherent light-matter interactions. For instance, whereas Floquet-Bloch states have been conventionally realized by hybridization between electronic states and an external periodic light field 79, 80, 81, 82, 83, 84, 85, there are a number of theoretical proposals for achieving the Floquet states by replacing photons with phonons 86, 87 or excitons 88. One signature of Floquet-Bloch states is the appearance of electronic band replicas detected in tr-ARPES experiments, where these replica bands follow specific intensity relations and polarization selection rules given by the exciting photons. Importantly, these bands are spectrally spaced apart by , where is the angular frequency of the bosonic field, be it a photon, phonon, or exciton. In a recent photoemission experiment using a momentum microscope specifically designed to study extremely small samples, exciton-induced Floquet effects were reported in monolayer WS2 in the presence of a high density of resonantly-driven photoexcited excitons 89, which also bear the hallmark of an excitonic insulator phase.
@@ II.4 Phase resolution of coherent collective excitations
When coherent dynamics are launched in solids upon photoexcitation, the motion of collective excitations manifests as wave-like oscillations. The waveform of these oscillations can be directly visualized via an ultrafast probe, an example of which is shown in Fig. 1b that depicts an oscillating Kerr rotation signal caused by a coherent magnon. Therefore, the phase of the collective excitation oscillations can be determined (labeled as in Fig. 1b, bottom panel), which equilibrium experiments cannot measure directly.
@@
Phase-sensitivity can be used to isolate a single collective mode that affects multiple degrees of freedom, providing a way to measure the coupling strength between them 19, 22. More specifically, by selectively inducing a particular oscillatory response using a tailored pump, and by leveraging phase-locked coherent features from different observables for the corresponding degrees of freedom, one can obtain the coupling strength directly, independent of the theoretical model. For example, following the excitation of a coherent phonon, tr-ARPES can probe spectral shifts in the electronic band energy while time-resolved X-ray diffraction (tr-XRD) can quantify the shifts in each atomic position for the phonon of interest. By combining these two measurements, electron-phonon coupling of a specific electronic band to a specific phonon mode can be calculated. This procedure has been applied to an iron-based superconductor FeSe (ref. 19) and three-dimensional topological insulators Bi2Se3 and Bi2Te3 (ref. 22). The extracted electron-phonon coupling strength in Bi2Se3 and Bi2Te3 is generally consistent with the expectation from density functional theory (DFT) computations while that in FeSe shows a significant deviation from DFT, which underestimates the strength. Instead, after electron-electron correlation effects were included in a self-consistent DFT-dynamical mean field theory, a quantitative agreement emerged, indicating an enhancement of the coupling strength due to strong electron correlations. Besides electron-phonon coupling, this multi-probe experimental protocol has also been deployed to measure the dynamical magnetoelectric coupling in multiferroic materials. For instance, in NiI2, a giant coupling strength was obtained 90 where a single electromagnon mode induces periodic modulations in the dipolar and magnetic orders, revealing an interesting /2 phase shift and lock-in behavior between electric polarization and magnetization oscillations that are at terahertz frequencies.
@@
Even with a single experimental probe, a change in the oscillation phase with respect to some tuning parameter can be used to reveal material properties 91, 92, 93 and identify microscopic processes underlying the collective excitation 94. For instance, far above the Curie temperature, the optical phonon mode in a layered ferromagnetic insulator CrSiTe3 can be excited due to the conventional displacive mechanism 15, but the coherent phonon phase experiences a shift when the material cools below a temperature scale where in-plane ferromagnetic correlations start to develop 92. Such a phase shift is only observed when the excitation photon energy is sufficiently large to induce charge transfer from the Te2- to Cr3+ ions, a process that also increases the superexchange energy. The phase shift is interpreted as evidence for a coherent phonon mechanism due to superexchange enhancement via magneto-elastic coupling, which only applies in the presence of short-range ferromagnetic correlation in this material. Notably, the phase shift of an mode across the Néel temperature also appears in an antiferromagnetic insulator FePS3, providing evidence for selective coupling between the magnetic order and the trigonal distortions in this layered system 93. Another example that highlights the importance of the oscillation phase comes from a coherent acoustic phonon in a freestanding film of La2/3Ca1/3MnO3. Three different observables in time-resolved diffraction (Bragg peak intensity, width, and position) were observed to synchronously oscillate with a particular phase relation — either in-phase or out-of-phase — after photoexcitation 94. By theoretically modeling the phase relation across the distinct observables from this coherent acoustic mode, Su and coworkers interpreted the phase relation as features of ultrafast wrinkle formation that results from delamination between the freestanding film and the substrate at its boundary, offering unique insights into the oft-neglected film–substrate interaction that governs the properties of freestanding structures.